1. Calculate the ionic strength of a 0.02 M solution of Al2(SO4)3


1
Expert's answer
2022-02-28T11:01:09-0500

[Al3+]=2×0.02 M

[SO42−​]=3×0.02 M

Ionic strength of solution =1/2 x ​[0.02×2×32+0.02×3×22] =0.3